OpenStudy (anonymous):

HELP Will Give Metal!!!! What is the value of y in the solution set of the system of linear equation shown below. *10 *8 *6 *2

OpenStudy (anonymous):

PLEASE HELP!!!! @NY,NY

OpenStudy (anonymous):

PLEASE HELP!!! @misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i would love to help, but he "below" is not shown!

OpenStudy (anonymous):

y=-x+12 4x-2y=36 Sorry @misty1212

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

lol hi again!

OpenStudy (misty1212):

you got \(y=-x+12\) which means the \(y\) in the second equation is also \(-x+12\)

OpenStudy (misty1212):

put it in and get \[4x-2(-x+12)=36\] and solve that sucker for \(x\) can u do that ?

OpenStudy (anonymous):

10? Can you possibly help me with 3 more? @misty1212

OpenStudy (misty1212):

hold on let me check your answer first

OpenStudy (misty1212):

yeah \(x=10\) but you still need \(y\) right? or no?

OpenStudy (misty1212):

actually that is all you need, it asks for \(y\) not \(x\) so we have to find \(y\)

OpenStudy (anonymous):

ok, so 10 is or isn't the answer? @misty1212

OpenStudy (misty1212):

no, 10 is x you are asked for y

OpenStudy (misty1212):

but you know \(y=-x+12\) so y is easy to find if \(x=10\)

OpenStudy (anonymous):

so 2? @misty1212

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

which of the following expressions is equivalent to x(x+5) - (1-5x)? *-4x^2+6x *x^2-1 *6x^2+4x *x^2+10-1 which inequality is equivalent to 5x-2y>10? *y>5/2x-10 *y> 5/2x+5 *y <2/5x - 5/2 *y< 5/2x-5 @misty1212

OpenStudy (misty1212):

\[x(x+5)-(1-5x)\\ x^2+5x-1+5x\\ x^2+10x-1\]

OpenStudy (misty1212):

\[5x-2y>10\\ 5x+10>2y\\ \frac{5}{2}x+5>y\]

OpenStudy (misty1212):

or \[y<\frac{5}{2}x+5\]

OpenStudy (anonymous):

I don't understand?? @misty1212

OpenStudy (misty1212):

where are you stuck?

OpenStudy (misty1212):

which one, first or second?

OpenStudy (anonymous):

I guess lets start with the first one...are those the steps to get the answer? @misty1212

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

so the first answer is D and the second answer is A??

OpenStudy (misty1212):

they are both D

OpenStudy (misty1212):

i made a typo, last one should be \[y<\frac{5}{2}x-5\]

OpenStudy (anonymous):

Thank you so much can you possibly help with one last question

OpenStudy (misty1212):

sure

OpenStudy (anonymous):

The perimeter of a rectangle is 90 centimeters. The width of the rectangle can be represented by (x+2), and its length can be represented by (3x-5)What are the dimensions of this rectangle in centimeters? *Width=10 Length 25 *Width=12 Length=29 *Width=14 Length=31 *Width=16 Length=32

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

\[2(x+2)+2(3x-5)=90\] solve for \(x\)

OpenStudy (anonymous):

12 @misty1212

OpenStudy (misty1212):

\[2x+4+6x-10=90\\ 8x-6=90\\ 8x=96\\ x=12\]

OpenStudy (misty1212):

yeah that is x but that is not the answer

OpenStudy (misty1212):

the dimensions are \(x+2\) which is \(14\) by \(3x-5\) which is 31

OpenStudy (anonymous):

so C